Express sin^5 x as a linear combination of sin kx and coskx, k belongs to Z. using De Moivre's theorem.
1
Expert's answer
2020-02-24T11:42:39-0500

Let's have a look on "\iota" complex number (Cos(x) + "\iota"*Sin(x)). And calculate the fifth power of this number

(Cos(x)+"\iota"*Sin(x))5 = Cos5(x) + 5"\iota"*Cos4(x)Sin(x) + 10"\iota"2Cos3(x)Sin2(x) + 10"\iota"3*Cos2(x)Sin3(x) + 5"\iota"4*Cos(x)Sin4(x) + "\iota"5Sin5(x) [using binomial formula] = (Cos5(x) - 10*Cos3(x)Sin2(x) + 5*Cos(x)Sin4(x)) + "\iota"*(5*Cos4(x)Sin(x) - 10*Cos2(x)Sin3(x) + Sin5(x)).

From the other side, (Cos(x)+"\iota"*Sin(x))5 = Cos(5*x) + "\iota"*Sin(5*x) [using Moivre's theorem]

It means that both numbers are equal.

So Sin(5x) = 5*Cos4(x)Sin(x) - 10*Cos2(x)Sin3(x) + Sin5(x) "\implies"

Sin5(x) = Sin(5x) - 5*Cos4(x)Sin(x) + 10*Cos2(x)Sin3(x) = Sin(5x) + 10*(1 - Sin2(x))Sin3(x) - 5*(1 - Sin2(x))2Sin(x) = Sin(5x) + 10*Sin3(x) - 10*Sin5(x) - 5*Sin(x) + 10*Sin3(x) - 5*Sin5(x) "\implies"

16*Sin5(x) = Sin(5x) + 20*Sin3(x) - 5*Sin(x)

/// [using the fornula for 4*Sin3(x) = 3*Sin(x) - Sin(3x)]

We get 16*Sin5(x) = Sin(5x) + 15*Sin(x) - 5*Sin(3x) - 5*Sin(x) = Sin(5x) - 5*Sin(3x) + 10*Sin(x)

So, the result is Sin5(x) = (1/16)*Sin(5x) - (5/16)*Sin(3x) + (5/8)*Sin(x)